OpenStudy (anonymous):

Solve sin(x + (π / 4)) – sin(x – (π / 4) = 1 (Be sure to show all of your work.) Will give medal & fan if someone walks me through this problem

OpenStudy (anonymous):

@whpalmer4 @jim_thompson5910 @Hero can any of you help?

OpenStudy (anonymous):

or you lol

myininaya (myininaya):

did you try using the sum and difference identity for sin?

myininaya (myininaya):

\[\sin(a+b)=\sin(a)\cos(b)+\sin(b)\cos(a)\]

OpenStudy (anonymous):

I tried but I just don't understand how that ties into the equation and whatnot

myininaya (myininaya):

well you have sin(x+pi/4) that is a sum inside so you can totally use the identity

myininaya (myininaya):

something might cancel you might make something ugly look a lot nicer

OpenStudy (whpalmer4):

I think you also need the corresponding \(\sin(a-b) = \sin(a)\cos(b) - \sin(b)\cos(a)\)

myininaya (myininaya):

yep I just wrote down one of them because you also just use just the sum identity

OpenStudy (whpalmer4):

although I guess it's 6 of 1, half dozen of the other

myininaya (myininaya):

along with knowing sin is odd function and cos is even function

OpenStudy (anonymous):

I just need everything explained to me step by step so I can write it own for future problems

myininaya (myininaya):

you said you tried to use those identities can i see what you got after using them?

OpenStudy (anonymous):

I didn't "try" to use them as much as look at the identities and the equation and become confuse

OpenStudy (anonymous):

*confused

OpenStudy (anonymous):

I really just want it in a step-by-step explanation because I don't understand much of it, and seeing it played out step by step is what helps me learn these things.

myininaya (myininaya):

I will work with you if you are willing to answer my questions as we go along with solving the problem. I will not just give you the answer though. I don't see how me just squirting out the answer would be of much help. I want you to tell me how to use the sum identity on sin(x+pi/4)

OpenStudy (whpalmer4):

it really is most useful to stumble around a few times while doing these, to get a feel for how to do it...

OpenStudy (whpalmer4):

and how not to do it :-)

OpenStudy (anonymous):

I didn't ask you to give me the straight up answer, but I'm just saying I'm already not going to understand anything you ask me.. I'll try to answer them the best I can though

myininaya (myininaya):

ok first question can you use the sum identity for sin on sin(x+pi/4) the identity is sin(a+b)=sin(a)cos(b)+sin(b)cos(a)

OpenStudy (anonymous):

yes..? I would guess so

myininaya (myininaya):

Ok will you do it

OpenStudy (anonymous):

That's the problem, first off: do what? and second: Whatever it is, I don't know how to do it /:

myininaya (myininaya):

 In this drawing if I wanted these two things to be equal what would a need to be and what would b need to be?